# Revenue Maximization Under Sequential Price Competition Via The Estimation Of $s$-Concave Demand Functions

### v3

@@ 3 Problem Formulation
$y_{i}^{(t)}=\lambda_{i}(\mathbf{p}^{(t)})+\varepsilon_{i}^{(t)},\quad\lambda_{i}(\mathbf{p}^{(t)})\triangleq\psi_{i}(-\beta_{i}p_{i}^{(t)}+\langle\bm{\gamma}_{i},\mathbf{p}_{-i}^{(t)}\rangle)=\psi_{i}(\langle\bm{\theta}_{i},\mathbf{p}^{(t)}\rangle),\quad t\in\mathcal{T},$ (6)
where $\{\varepsilon_{i}^{(t)}\}_{t\in\mathcal{T}}$ are (zero mean) $\sigma_{i}$-sub-gaussian demand noises following independent and identical distributions ($\varepsilon_{i}^{(t)}$ and $\varepsilon_{j}^{(t)}$ can be correlated with $i\neq j$, $i,j\in\mathcal{N}$), and $\bm{\theta}_{i}$ is an unknown vector of dimension $N$ with $i$-th entry equal to $\beta_{i}$ and the remaining entries being the values of $\bm{\gamma}_{i}\in\mathbb{R}^{N-1}$, ordered. The parameter vector $\bm{\gamma}_{i}$ measures how seller $i$ ’s demand is affected by competitor prices. We assume that the parameter space is such that the average demand $\lambda_{i}$ is non-negative and $\partial_{p_{i}}\lambda_{i}<0$ among all values of $\{\bm{\theta}_{i},\psi_{i}\}_{i\in\mathcal{N}}$; a similar assumption is found in Birge et al. (2024); Li et al. (2024). The above conditions hold if $\psi_{i},\psi_{i}^{\prime}>0$ and $\beta_{i}>0$, which are explicitly assumed in ˜3.1.
@@
The constants $\underline{B}_{\psi_{i}}$ and $\bar{B}_{\psi_{i}}$ are known by the seller $i$ (they are used to compute the optimization problem in (11)). The condition $\|\bm{\theta}_{i}\|_{2}=1$ guarantees the differentiability of the monotone single index model (Chmielewski, 1981; Balabdaoui et al., 2019). Note that Equation˜6 is well defined, that is $\langle\bm{\theta}_{i},\mathbf{p}^{(t)}\rangle\in\mathcal{U}$, indeed $\|\bm{\theta}_{i}\|_{2}=1\implies|\langle\bm{\theta}_{i},\mathbf{p}\rangle|\leq\|\mathbf{p}\|_{2}\leq p_{\max}$, $\forall\mathbf{p}\in\mathcal{P}$.
3.1 Individual Regret
@@
$\mathrm{R}_{i}(T)\triangleq\mathbb{E}\sum_{t=1}^{T}p_{i}^{(t)}y_{i}^{(t)}=\mathbb{E}\sum_{t=1}^{T}\operatorname{rev}_{i}(p_{i}^{(t)}\mid\mathbf{p}^{(t)}_{-i}),\quad\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i})\triangleq p_{i}\psi_{i}(-\beta_{i}p_{i}+\langle\bm{\gamma}_{i},\mathbf{p}_{-i}\rangle).$
@@
$\operatorname{Reg}_{i}(T)\triangleq\left[\mathbb{E}\sum_{t=1}^{T}\operatorname{rev}_{i}(\Gamma_{i}(\mathbf{p}^{(t)}_{-i})\mid\mathbf{p}^{(t)}_{-i})\right]-\mathrm{R}_{i}(T),\text{ }\Gamma_{i}(\mathbf{p}_{-i})\in\underset{p_{i}\in\mathcal{P}_{i}}{\operatorname{argmax}}\,\operatorname{rev}_{i}(p_{i}\mid\mathbf{p}_{-i}),$ (7)
@@ 3.2 Nash Equilibrium (NE)
where $\bm{\Gamma}$ is called Best Response Operator, and its coponents $\Gamma_{i}$ are defined in (7).
Before determining the Nash equilibrium (NE), we emphasize why its existence and uniqueness is crucial, and what conditions guarantee convergence to the NE.
The objective of this paper is to design an algorithm that guarantees sublinear regret for every seller, i.e., $\operatorname{Reg}_{i}(T)=o(T)$ for every $i\in\mathcal{N}$. Our policy consists of two phases: an exploration phase, in which each seller learns their individual best-response map $\hat{\Gamma}_{i}$ by consistently estimating the parameters $(\bm{\theta}_{i},\psi_{i})$, and an exploitation phase, in which, at each round $t$, seller $i$ sets their price according to $p_{i}^{(t)}=\hat{\Gamma}_{i}(\mathbf{p}_{-i}^{(t-1)})$, or equivalently $\mathbf{p}^{(t)}=\hat{\bm{\Gamma}}(\mathbf{p}^{(t-1)})$. The link between regret and equilibrium is straightforward: the individual regret is controlled if the iterates converge to the NE. In fact, $\operatorname{Reg}_{i}(T)\lesssim T\mathbb{E}\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}$ (see (37) for a detailed bound). Now, assume that there is a unique $\mathbf{p}^{\star}$. By the triangle inequality, we have
@@
where $L_{\bm{\Gamma}}$ is the Lipschitz constant of $\bm{\Gamma}$ and $\|\mathbf{F}\|_{\infty}=\sup_{\mathbf{p}\in\mathcal{P}}\|\mathbf{F}(\mathbf{p})\|$. Therefore, convergence to $\mathbf{p}^{\star}$ requires two ingredients:
  1. 1.
    Contraction: $0<L_{\bm{\Gamma}}<1$ which ensures that deviations from the NE shrink geometrically;
  2. 2.
    Consistent estimation for $\|\hat{\bm{\Gamma}}-\bm{\Gamma}\|_{\infty}$, which follows from consistent estimation of $(\bm{\theta}_{i},\psi_{i})$ and strong concavity of $\operatorname{rev}_{i}(\cdot\mid\mathbf{p}_{-i})$ uniformly in $\mathbf{p}_{-i}$, for each seller $i\in\mathcal{N}$.
Putting these together, we obtain (see (G.2) for a formal inequality):
@@
The uniqueness of the Nash equilibrium is essential for regret convergence. If two distinct NEs $\mathbf{p}^{\star}\neq\mathbf{q}^{\star}$ exist, then depending on initialization and estimation error, different coordinates of $\mathbf{p}^{(t)}$ may converge to different equilibria (e.g., $p_{i}^{(t)}\to p_{i}^{\star}$ while $p_{j}^{(t)}\to q_{j}^{\star}$). In this case, $\|\mathbf{p}^{(t)}-\mathbf{p}^{\star}\|$ (or $\|\mathbf{p}^{(t)}-\mathbf{q}^{\star}\|$) would not vanish, so the bound $\operatorname{Reg}_{i}(T)\lesssim\mathbb{E}\|\mathbf{p}^{(T)}-\mathbf{p}^{\star}\|^{2}\to 0$ would fail, and regret convergence would not be guaranteed.
Both strong concavity and the contraction property independently suffice to guarantee the uniqueness of the NE. However, as we have discussed, for the convergence of the algorithm, both are required.
In what follows, we proceed in the following steps: in Step 1 we establish existence and uniqueness of the NE using strong concavity of the revenue functions; in Step 2 we derive an explicit representation of the NE, based on shape constraints of the functions $\psi_{i}$, that will be used to estimate $\Gamma_{i}$; we conclude with Step 3, by imposing conditions guaranteeing contraction of $\bm{\Gamma}$.
Step 1: Existence and Uniqueness of the NE.
@@
By twice differentiability of $\psi_{i}$, ˜3.2 is guaranteed if $\mu_{i}\triangleq 2\beta_{i}\underline{B}_{\psi_{i}^{\prime}}-\beta_{i}^{2}\bar{p}_{i}B_{\psi_{i}^{\prime\prime}}>0$. Indeed, since by ˜3.1 $|\psi_{i}^{\prime\prime}|\leq B_{\psi_{i}^{\prime\prime}}$ and $\psi_{i}^{\prime}\geq B_{\psi^{\prime}_{i}}$, we have
@@
˜3.2 guarantees that the best-response map $\Gamma_{i}$ in (7) is well defined for every $\mathbf{p}_{-i}$. More importantly, it guarantees the uniqueness of the Nash equilibrium, in line with standard results in competitive games (Scutari et al., 2014; Li et al., 2024; Tsekrekos & Yannacopoulos, 2024). Instances where ˜3.2 is satisfied include linear demand models (Li et al., 2024), concave demand specifications, and a class of $s_{i}$-concave demand functions with $s_{i}>-1$. A more detailed discussion of these and further examples is provided in Section˜I.1. The next result is an immediate consequence of Theorem 3, part (d), in Scutari et al. (2014).
Lemma 3.3 (Existence and Uniqueness of NE).
Under assumptions (3.1) and (3.2), there exists a unique $\mathbf{p}^{*}\in\mathcal{P}$ satisfying the fixed point equation in (8).
Step 2: $s$-concavity implies the First Order Conditions.
Under ˜3.2, the map $\Gamma_{i}(\mathbf{p}_{-i})$ can be recovered by solving the first order conditions (FOCs), $\partial_{p_{i}}\operatorname{rev}_{i}(p_{i}\mid\mathbf{p}_{-i})=0$, and projecting the solution onto $\mathcal{P}_{i}$. We now determine a shape constraint assumption that is sufficient to satisfy the FOCs. Specifically, we can show that $\Gamma_{i}$ can be written as (see the proof of Lemma˜3.7 for the derivation)
$\Gamma_{i}(\mathbf{p}_{-i})=\Pi_{\mathcal{P}_{i}}\nicefrac{{g_{i}(\langle\bm{\gamma}_{i},\mathbf{p}_{-i}\rangle)}}{{\beta_{i}}},\quad g_{i}(u)\triangleq u-\varphi_{i}^{-1}(u),\quad\varphi_{i}(u)\triangleq u+\frac{\psi_{i}(u)}{\psi_{i}^{\prime}(u)},$
provided $\varphi_{i}^{-1}$ exists. Here $\Pi_{\mathcal{P}_{i}}$ is the projection into $\mathcal{P}_{i}$ and $\varphi_{i}$ is called virtual valuation function of firm $i$. For this purpose, we introduce ˜3.4, which makes the $\varphi_{i}$’s invertible.
@@
For every $i\in\mathcal{N}$, there exists a constant $c_{i}>0$, known to seller $i$, such that $\varphi_{i}^{\prime}\geq c_{i}$.
˜3.4 covers the linear demand model by Li et al. (2024) for which $\psi_{i}(u)=\alpha_{i}+u\implies\varphi_{i}(u)=2u+\alpha_{i}\implies\psi_{i}^{\prime}\geq 2$ and can also be found in several works with $N=1$ (monopolistic setting); for example, Assumption 2.1 in Fan et al. (2024), Assumption 1 in Chen & Farias (2018), Equation 1 in Cole & Roughgarden (2014) and Golrezaei et al. (2019); Javanmard (2017); Javanmard & Nazerzadeh (2019) which assume $\psi_{i}$ and $1-\psi_{i}$ to be log-concave (which specifically implies $c_{i}\geq 1$), where $\psi_{i}$ is a survival function.
We now present an equivalent formulation of ˜3.4 in terms of $s$-concavity, a condition that will allow us to estimate $\psi_{i}$ via shape constraints. The proof is provided in Appendix˜B.
@@
$\sup_{i\in\mathcal{N}}\|g_{i}^{\prime}\|_{\infty}\frac{\|\bm{\gamma}_{i}\|_{1}}{\beta_{i}}<1$.
˜3.6 generalizes Assumption 2 in Li et al. (2024) (in their case $\psi_{i}(u)=u+\alpha_{i}\implies g_{i}(u)=(u+\alpha_{i})/2\implies g_{i}^{\prime}=1/2$, matching the contraction constant $L_{\bm{\Gamma}}$ in Equation 30 of their work). Similar assumptions are found in Kachani et al. (2007). The proof of the following result is relegated to Appendix˜C.
@@
Let Assumptions 3.2, 3.4 and 3.6 hold. Then, $\bm{\Gamma}=(\Gamma_{i})_{i\in[N]}:\mathcal{P}\to\mathcal{P}$ is a contraction with contraction constant $L_{\bm{\Gamma}}:=\sup_{i\in\mathcal{N}}\|g_{i}^{\prime}\|_{\infty}\nicefrac{{\|\bm{\gamma}_{i}\|_{1}}}{{\beta_{i}}}<1$.
@@
$g_{i}^{\prime}(u)=1-\frac{1}{\varphi_{i}^{\prime}(\varphi_{i}^{-1}(u))}\geq 1-\frac{1}{c_{i}}>-1,\quad g_{i}^{\prime}(u)=1-\frac{1}{\varphi_{i}^{\prime}(\varphi_{i}^{-1}(u))}<1,\quad\Rightarrow\|g_{i}^{\prime}\|_{\infty}<1.$
In this case ˜3.6 reduces to $\beta_{i}>\|\bm{\gamma}_{i}\|_{1}$, but since $\beta_{i}=(1-\|\bm{\gamma}_{i}\|_{2}^{2})^{1/2}$, the condition becomes $\|\bm{\gamma}_{i}\|_{1}^{2}+\|\bm{\gamma}_{i}\|_{2}^{2}<1$. But since $\|\bm{\gamma}_{i}\|_{2}\leq\|\bm{\gamma}_{i}\|_{1}$, a sufficient condition is $\|\bm{\gamma}_{i}\|_{1}<\nicefrac{{1}}{{\sqrt{2}}}$.

### v4

@@ 3 Problem Formulation
$\textstyle y_{i}^{(t)}=\lambda_{i}(\mathbf{p}^{(t)})+\varepsilon_{i}^{(t)},\quad\lambda_{i}(\mathbf{p}^{(t)})\triangleq\psi_{i}(-\beta_{i}p_{i}^{(t)}+\langle\bm{\gamma}_{i},\mathbf{p}_{-i}^{(t)}\rangle)=\psi_{i}(\langle\bm{\theta}_{i},\mathbf{p}^{(t)}\rangle),\quad t\in\mathcal{T},$ (6)
where $\{\varepsilon_{i}^{(t)}\}_{t\in\mathcal{T}}$ are (zero mean) $\sigma_{i}$-sub-gaussian demand noises following independent and identical distributions ($\varepsilon_{i}^{(t)}$ and $\varepsilon_{j}^{(t)}$ can be correlated with $i\neq j$, $i,j\in\mathcal{N}$), and $\bm{\theta}_{i}$ is an unknown vector of dimension $N$ with $i$-th entry equal to $\beta_{i}$ and the remaining entries being the values of $\bm{\gamma}_{i}\in\mathbb{R}^{N-1}$, ordered. The parameter vector $\bm{\gamma}_{i}$ measures how seller $i$ ’s demand is affected by competitor prices. We assume that the parameter space is such that the average demand $\lambda_{i}$ is non-negative and $\partial_{p_{i}}\lambda_{i}<0$ among all values of $\{\bm{\theta}_{i},\psi_{i}\}_{i\in\mathcal{N}}$; a similar assumption is found in birge2024interfere; li2024lego. The above conditions hold if $\psi_{i},\psi_{i}^{\prime}>0$ and $\beta_{i}>0$, which are explicitly assumed in 3.1.
@@
The constants $\underline{B}_{\psi_{i}},\bar{B}_{\psi_{i}}$ are known by the seller $i$ (they are used to compute the optimization problem in (11)), however, firms can estimate them easily from historical sales data and operational capacity limits; moreover loose bounds suffice and do not affect the rates in Theorem 5.4. The condition $\|\bm{\theta}_{i}\|_{2}=1$ guarantees the differentiability of the monotone single index model (chmielewski1981elliptically; balabdaoui2019least). Note that Equation 6 is well defined, that is $\langle\bm{\theta}_{i},\mathbf{p}^{(t)}\rangle\in\mathcal{U}$, indeed $\|\bm{\theta}_{i}\|_{2}=1\implies|\langle\bm{\theta}_{i},\mathbf{p}\rangle|\leq\|\mathbf{p}\|_{2}\leq p_{\max}$, $\forall\mathbf{p}\in\mathcal{P}$. Regarding $\bm{\gamma}_{i}$, while many applications assume $\gamma_{ij}>0$, this restriction is not required for our algorithm or theoretical guarantees - our results remain valid when $\gamma_{ij}>0$. We allow $\gamma_{ij}$ to be negative to capture settings with negative cross-effects, such as vertically differentiated products, where raising a competitor’s price may reduce my demand if consumers perceive their product as higher quality.
Individual Regret.
@@
$\textstyle\mathrm{R}_{i}(T)\triangleq\mathbb{E}\sum_{t=1}^{T}p_{i}^{(t)}y_{i}^{(t)}=\mathbb{E}\sum_{t=1}^{T}\operatorname{rev}_{i}(p_{i}^{(t)}\mid\mathbf{p}^{(t)}_{-i}),\quad\operatorname{rev}_{i}(p_{i}|\mathbf{p}_{-i})\triangleq p_{i}\psi_{i}(-\beta_{i}p_{i}+\langle\bm{\gamma}_{i},\mathbf{p}_{-i}\rangle).$
@@
$\textstyle\operatorname{Reg}_{i}(T)\triangleq\left[\mathbb{E}\sum_{t=1}^{T}\operatorname{rev}_{i}(\Gamma_{i}(\mathbf{p}^{(t)}_{-i})\mid\mathbf{p}^{(t)}_{-i})\right]-\mathrm{R}_{i}(T),\text{ }\Gamma_{i}(\mathbf{p}_{-i})\in\underset{p_{i}\in\mathcal{P}_{i}}{\operatorname{argmax}}\,\operatorname{rev}_{i}(p_{i}\mid\mathbf{p}_{-i}),$ (7)
@@
where $\bm{\Gamma}$ is called Best Response Operator, and its components $\Gamma_{i}$ are defined in (7).
3.1 Main Assumptions
Before presenting the main assumptions, we clarify why our framework requires two key properties: the best-response map must be contractive, and each seller’s revenue function must be strongly concave. These conditions ensure well-posedness of the equilibrium and enable the stability and convergence guarantees developed later.
The objective of this paper is to design an algorithm that guarantees sublinear regret for every seller, i.e., $\operatorname{Reg}_{i}(T)=o(T)$ for every $i\in\mathcal{N}$. Our policy consists of two phases: an exploration phase, in which each seller learns their individual best-response map $\hat{\Gamma}_{i}$ by consistently estimating the parameters $(\bm{\theta}_{i},\psi_{i})$, and an exploitation phase, in which, at each round $t$, seller $i$ sets their price according to $p_{i}^{(t)}=\hat{\Gamma}_{i}(\mathbf{p}_{-i}^{(t-1)})$, or equivalently $\mathbf{p}^{(t)}=\hat{\bm{\Gamma}}(\mathbf{p}^{(t-1)})$. The link between regret and equilibrium is straightforward: the individual regret is controlled if the iterates converge to the NE. In fact, $\operatorname{Reg}_{i}(T)\lesssim T\mathbb{E}\|\mathbf{p}^{(T)}-\mathbf{p}^{*}\|_{2}^{2}$, were $\mathbf{p}^{\star}$ is a NE (see (37) for a detailed bound). By the triangle inequality, we have
@@ 3.1 Main Assumptions
where $L_{\bm{\Gamma}}$ is the Lipschitz constant of $\bm{\Gamma}$ and $\|\mathbf{F}\|_{\infty}=\sup_{\mathbf{p}\in\mathcal{P}}\|\mathbf{F}(\mathbf{p})\|$. Therefore, convergence to $\mathbf{p}^{\star}$ requires two ingredients: $(i)$ Contraction: $0<L_{\bm{\Gamma}}<1$ which ensures that deviations from the NE shrink geometrically; $(ii)$ Consistent estimation for $\|\hat{\bm{\Gamma}}-\bm{\Gamma}\|_{\infty}$, which follows from consistent estimation of $(\bm{\theta}_{i},\psi_{i})$ and strong concavity of $\operatorname{rev}_{i}(\cdot\mid\mathbf{p}_{-i})$ uniformly in $\mathbf{p}_{-i}$, for each seller $i\in\mathcal{N}$. Putting these together, we obtain (see (G.2) for a formal inequality):
@@
Additionally, strong concavity ensures the existence of the NE, while the contraction property of the best response map implies the uniqueness of the NE.
@@
By twice differentiability of $\psi_{i}$, 3.2 is guaranteed if $\mu_{i}\triangleq 2\beta_{i}\underline{B}_{\psi_{i}^{\prime}}-\beta_{i}^{2}\bar{p}_{i}B_{\psi_{i}^{\prime\prime}}>0$. Indeed, since by 3.1 $|\psi_{i}^{\prime\prime}|\leq B_{\psi_{i}^{\prime\prime}}$ and $\psi_{i}^{\prime}\geq B_{\psi^{\prime}_{i}}$, we have
@@
3.2 guarantees that the best-response map $\Gamma_{i}$ in (7) is well defined for every $\mathbf{p}_{-i}$. More importantly, it guarantees the existence of the Nash equilibrium, in line with standard results in competitive games (scutari2014real; li2024lego; tsekrekos2024variational). Instances where 3.2 is satisfied include linear demand models (li2024lego), concave demand specifications, and a class of $s_{i}$-concave demand functions with $s_{i}>-1$. A more detailed discussion of these and further examples is provided in Section I.1. The next result is an immediate consequence of Theorem 3, in scutari2014real.
Lemma 3.3 (Existence of NE).
Under assumptions (3.1) and (3.2), there exists a $\mathbf{p}^{*}\in\mathcal{P}$ satisfying the fixed point equation in (8).
By 3.2, the map $\Gamma_{i}(\mathbf{p}_{-i})$ can be recovered by solving the first order conditions (FOCs), $\partial_{p_{i}}\operatorname{rev}_{i}(p_{i}\mid\mathbf{p}_{-i})=0$, and projecting the solution onto $\mathcal{P}_{i}$. We now determine a shape constraint assumption that is sufficient to satisfy the FOCs. Specifically, we can show that $\Gamma_{i}$ can be written as (see the proof of Lemma 3.7 for the derivation)
$\Gamma_{i}(\mathbf{p}_{-i})=\Pi_{\mathcal{P}_{i}}\nicefrac{{g_{i}(\langle\bm{\gamma}_{i},\mathbf{p}_{-i}\rangle)}}{{\beta_{i}}},\quad g_{i}(u)\triangleq u-\varphi_{i}^{-1}(u),\quad\varphi_{i}(u)\triangleq u+\nicefrac{{\psi_{i}(u)}}{{\psi_{i}^{\prime}(u)}},$
provided $\varphi_{i}^{-1}$ exists. Here $\Pi_{\mathcal{P}_{i}}$ is the projection into $\mathcal{P}_{i}$ and $\varphi_{i}$ is called virtual valuation function of firm $i$. For this purpose, we introduce 3.4, which makes the $\varphi_{i}$’s invertible.
@@
$\forall i\in\mathcal{N}$, there exists a constant $c_{i}>0$, known to seller $i$, such that $\varphi_{i}^{\prime}\geq c_{i}$.
3.4 covers the linear demand model by li2024lego for which $\psi_{i}(u)=\alpha_{i}+u\implies\varphi_{i}(u)=2u+\alpha_{i}\implies\psi_{i}^{\prime}\geq 2$ and can also be found in several works with $N=1$ (monopolistic setting); for example, Assumption 2.1 in fan2024policy, Assumption 1 in chen2018robust, Equation 1 in cole2014sample and golrezaei2019dynamic; javanmard2017perishability; javanmard2019dynamic which assume $\psi_{i}$ and $1-\psi_{i}$ to be log-concave (which specifically implies $c_{i}\geq 1$), where $\psi_{i}$ is a survival function.
We now present an equivalent formulation of 3.4 in terms of $s$-concavity, a condition that will allow us to estimate $\psi_{i}$ via shape constraints (the proof is provided in Appendix B). Examples of $s$-concave functions can be found in Appendix L.
@@
$\sup_{i\in\mathcal{N}}\|g_{i}^{\prime}\|_{\infty}\nicefrac{{\|\bm{\gamma}_{i}\|_{1}}}{{\beta_{i}}}<1$.
3.6 generalizes Assumption 2 in li2024lego (in their case $\psi_{i}(u)=u+\alpha_{i}\implies g_{i}(u)=(u+\alpha_{i})/2\implies g_{i}^{\prime}=1/2$, matching the contraction constant $L_{\bm{\Gamma}}$ in Equation 30 of their work). Similar assumptions are found in kachani2007modeling. The proof of the following result is relegated to Appendix C. In practice, 3.6 states that the influence of competitors’ prices on seller $i$’s optimal response is sufficiently small relative to the sensitivity to its own price.
@@
Let Assumptions 3.2, 3.4 and 3.6 hold. Then, $\bm{\Gamma}=(\Gamma_{i})_{i\in[N]}:\mathcal{P}\to\mathcal{P}$ is a contraction with contraction constant $L_{\bm{\Gamma}}:=\sup_{i\in\mathcal{N}}\|g_{i}^{\prime}\|_{\infty}\nicefrac{{\|\bm{\gamma}_{i}\|_{1}}}{{\beta_{i}}}<1$. Consequently, the NE $\mathbf{p}^{*}$ is unique.
@@
$g_{i}^{\prime}(u)=1-\nicefrac{{1}}{{\varphi_{i}^{\prime}(\varphi_{i}^{-1}(u))}}\geq 1-\nicefrac{{1}}{{c_{i}}}>-1,\quad g_{i}^{\prime}(u)=1-\nicefrac{{1}}{{\varphi_{i}^{\prime}(\varphi_{i}^{-1}(u))}}<1,\quad\Rightarrow\|g_{i}^{\prime}\|_{\infty}<1.$
In this case 3.6 reduces to $\beta_{i}>\|\bm{\gamma}_{i}\|_{1}$, but since $\beta_{i}=(1-\|\bm{\gamma}_{i}\|_{2}^{2})^{1/2}$, the condition becomes $\|\bm{\gamma}_{i}\|_{1}^{2}+\|\bm{\gamma}_{i}\|_{2}^{2}<1$. But since $\|\bm{\gamma}_{i}\|_{2}\leq\|\bm{\gamma}_{i}\|_{1}$, a sufficient condition is $\|\bm{\gamma}_{i}\|_{1}<\nicefrac{{1}}{{\sqrt{2}}}$.